• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dhamilton

    Log entry time 08:17:32 on April 04, 2008

    Entry number 229857

    keyword=Owl Shift Summary

    Shift started at the end of BB MWDC HV/threshold scan.
    Run 2277 was the last setting for this sca (1450V, 12.5V).

    00:20 - Start of elastic optics data-taking.
    Sieve slit moved into position.
    BB MWDC HV set to 1450V.
    BB MWDC threshold set to 5V.
    BB coil current still at 390A.
    Run 2278 (BB 390A, 100k events) - ended early due to loss of beam.

    01:10 - MCC called to say that beam will be off while they cycle their
    magnets.

    02:20 - MCC called again. There is a problem with one of the arcs.
    Someone has been called to come in and try to fix it.

    05:00 - beam restored to hall.
    Run 2279 (BB 390A) - ended early.
    Immediate BB HV trip due to bad beam alignment.
    Target changed to BeO and alignement checked.

    05:55 - restart data-taking.
    Another HV trip due to drift in beam y position.

    05:30 - target changed back to 6cm LH2. Data-taking resumes.
    Run 2280 (BB 390A) - ended early.
    Another HV trip due to instability in beam y position.

    06:00 - move to empty target and request harp scan (all sigmas less than
    200 microns).

    06:45 - move to BeO target to check spot position again (still fine).

    06:55 - move to LH2 target to restart data-taking.
    Run 2281 (BB 390A) - ended early.
    The same problem in beam position instability.

    07:20 - spoke to the crew chief.
    They are going to try to identify the source of the problem. He suggested
    leaving the running conditions unchanged for now.

    07:30 - MCC have had some success identifying that the problem appears to
    be occurring before the locks.
    Changing to empty target to allow them to investigate further.

    07:50 - BB MWDC HV switched off.

    08:00 - shift ended with no beam and no solution to the problem with
    periodic instability in beam y position.